OpenStudy (anonymous):

Factor x^7 + 20x^5 + 100x^3 completely

OpenStudy (anonymous):

take out x^3, so it would be x^3(x^4+20x^2+100) now, let x^2 = a then you have a^3/2( a^2+20a+100) = a^3/2( a+10)^2 substituting a = x^2, x^3(x^2+10)^2